# Tally: Non-Intrusive Performance Isolation for Concurrent Deep Learning Workloads

### v1

@@ 3. GPU SHARING IN THE WILD
High integration and maintenance cost. Many existing GPU sharing systems are fundamentally intrusive in nature, necessitating extensive modifications to user code to be used in practice. This is particularly evident in application-level systems, where users must adapt their workloads to the specific programming interfaces proposed by these solutions. Such intrusive approaches not only burden the users, but also limit their flexibility in application development (lucid, ). Similarly, runtime-level GPU sharing solutions that are designed as extensions of DL frameworks impose a significant burden on cluster managers to implement and maintain them across diverse frameworks used by the users. Consequently, deploying these solutions in practice often involves substantial integration and maintenance costs, especially given the rapidly evolving landscape of the modern DL ecosystem.
@@
On the other hand, runtime-level GPU sharing solutions that seamlessly integrate with the container orchestration system or GPU driver, although may seem non-intrusive, they often impose implicit requirements on user code to gather additional application context needed for making effective scheduling decisions. This include requirements such as enabling the ability to temporarily launch applications in a specialized profiling environment before execution (orion2024, ), or mandating applications to implement checkpoint and restore mechanisms (gavel2020, ; xiao2018, ). Enforcing such programming practices in large-scale multi-tenant clusters is often impractical (Weng2022mlaas, ).
Lack of performance isolation guarantees. In production DL clusters, many inference tasks are governed by service-level agreements (SLAs) which impose strict latency requirements on their response times, typically of the order of a few milliseconds to seconds (clipper2017, ). However, current GPU sharing solutions often lack support for performance isolation, resulting in unacceptable performance degradation for latency-critical inference tasks when co-located with other tasks (GILMAN2021102234, ). For instance, NVIDIA’s native runtime-level GPU sharing solution, MPS (nvidia-mps, ), is designed to eagerly schedule as many GPU kernels from co-located tasks as possible. Although this design maximizes the GPU utilization, it can cause significant queuing delays for executing kernels of high-priority inference tasks. In such scenarios, our experiments (see Section 5) indicate that, the tail latency of a co-located inference task can go as high as 20×20$\times$20 × its standalone performance.
Effectively sharing GPU resources while ensuring performance isolation for latency-critical tasks requires the ability to prioritize one task over the other and rapidly switch between them based on their resource demands. The speed of context-switching between tasks, which we denote as turnaround latency, is dictated by the scheduling granularity used in GPU sharing systems. As described in Section 2, the scheduling granularity used by prior works primarily fall into four categories, from coarsest to finest: iteration-level (e.g. Salus (salus2020, )), kernel-level (e.g., TGS (tgs2023, )), block-level (e.g. EffiSha (effisha2017, )), and thread-level (e.g. REEF (reef2022, )). To illustrate the effect of scheduling granularity on the the tail latency, we present a representative empirical study in Table 1 where a latency-critical BERT (devlin-etal-2019-bert, ) inference task is co-executed with a Whisper (radford2023robust, ) model training workload on an NVIDIA A100 GPU. In this experiment, we empirically estimate the turnaround latency for different scheduling granularities on the Whisper model and compare it to the response time of BERT inference. Since BERT inference takes only $3.93$ ms, providing strong performance isolation of such tasks require sub-millisecond-level turnaround latency. However, both iteration-level and kernel-level scheduling can only achieve turnaround latencies of $3$ s and $10$ ms, respectively. As a result, scheduling the Whisper training task at these granularities can introduce 2.5760×2.5-760$\times$2.5 - 760 × higher overhead on BERT’s latency. In contrast, block-level and thread-level scheduling achieve microsecond-scale turnaround latencies of $304$ µs and $38$ µs respectively resulting in only mere $1-7\%$ overhead.
From these results, we conclude that to achieve practical performance isolation guarantees, GPU sharing solutions must schedule resources at either the block-level or thread-level granularity. Unfortunately, existing solutions either do not meet these criteria (e.g., AntMan, TGS) or prove ineffective in diverse production DL clusters due to their intrusive nature (e.g., EffiSha, Orion) or lack of generalizability (e.g. REEF). This brings us to the third limitation.
@@
Reliance on narrowly applicable workload characteristics. Many state-of-the-art GPU sharing systems rely on specific workload characteristics in their scheduling strategies. For instance, AntMan and TGS can provide runtime-level GPU sharing support but are limited to training workloads with predictable GPU kernel execution patterns. REEF, on the other hand, offers strong performance isolation for inference workloads, however, needs the GPU kernels to be idempotent to support thread-level kernel scheduling strategy, i.e., the kernel must always produce the same output for the same input, regardless of how many times it is executed. While these systems demonstrate benefits for certain class of applications, their underlying assumptions are often not applicable to large-scale multi-tenant DL clusters that must support a diverse range of workloads with varying characteristics. Prior studies (Weng2022mlaas, ) have revealed that DL clusters are often agnostic to the specific semantics of running workloads and extracting application-level context is largely infeasible. As a result, GPU sharing solutions that depend on specific task characteristics are not viable in many DL clusters.
@@
In summary, we argue that a GPU sharing solution for large-scale DL clusters must address all three of the aforementioned limitations.
@@ 4. TALLY: SYSTEM DESIGN
To facilitate the practical deployment of GPU sharing in production clusters, we propose Tally, a novel non-intrusive GPU sharing system that offers robust performance isolation for latency-critical tasks during shared execution with throughput-oriented best-effort tasks on a diverse range of DL workloads. Tally is designed as a virtualization layer operating transparently between applications and GPUs, intercepting device API calls and employs a task-agnostic scheduling algorithm to prioritize execution from high-priority tasks while opportunistically scheduling kernels from best-effort tasks during GPU idle cycles.
@@
Tally ensures performance isolation for high-priority tasks by scheduling kernels from best-effort tasks at the block-level granularity using two scheduling primitives: slicing and preemption. Slicing divides a large kernel into numerous smaller segments, enabling more controlled scheduling, whereas preemption allows for the interruption of active kernels. Importantly, we make two key observations that allows applying these two scheduling primitives in a task-agnostic and non-intrusive manner. First, unlike thread-level scheduling, thread blocks within a GPU are guaranteed to operate independently, making block-level scheduling universally applicable across GPU kernels. Moreover, while thread-level scheduling can provide stronger performance isolation, we find that block-level scheduling is sufficient for most practical scenarios. Second, the slicing and preemption scheduling primitives can be applied on GPU kernels through a series of transformation passes on the kernel device code (PTX in case of NVIDIA GPUs) obtained through the interception of device API calls. Leveraging these insights, we integrate the slicing and preemption primitives to Tally’s scheduling algorithm enabling microsecond-scale turnaround latency, allowing Tally to promptly switch execution between high-priority and best-effort tasks and maintain the millisecond-scale latency requirements in many latency-critical tasks.
Figure 1. Tally architecture
Figure 1 shows the high-level design of Tally, following a client-server architecture. On the application side, Tally’s client library intercepts device code and API calls initiated by client processes and forwards them to the Tally server, which is responsible for managing the actual device execution and scheduling. The objective of the Tally server is to facilitate priority-enforced concurrent execution of one high-priority task alongside multiple best-effort tasks. To achieve this, Tally server incorporates three key components: a (1) kernel transformer, which leverages intercepted device code to convert kernels into slicing and preemptive-style kernels, a (2) transparent profiler, which profiles the performance of the transformed kernels under various launch configurations at runtime and predicts their estimated turnaround latencies, and finally a (3) priority-aware scheduler, which integrates capabilities of the aforementioned modules and strategically schedules GPU execution of concurrent processes in a priority-aware manner. In the following sections, we provide details regarding each of these components.
@@
Figure 2. Kernel Transformations: (a) Slicing and Preemption Transformation (b) Unified Synchronization Transformation
@@ 4.1. Kernel Transformations
Slicing transformation. Slicing is a technique that facilitates fine-grained control over kernel execution by partitioning the thread blocks of a kernel into multiple sub-kernels. Figure 2(a) illustrates this concept, where a kernel originally composed of $16$ thread blocks is divided into $4$ sub-kernels, each containing $4$ blocks. As thread blocks within a kernel are guaranteed independent, theoretically, this partitioning should preserve the original kernel’s functional semantics. However, naively launching the original kernel $4$ times with $4$ thread blocks each will yield incorrect results, as GPU threads rely on built-in variables such as blockIdx to determine their task assignment. Specifically, the blockIdx variable informs a thread of the index of the block it resides in within the kernel execution, allowing it to properly compute its assigned task index. In the original kernel, threads are exposed to blockIdx values ranging from 00 to $15$ (as $16$ blocks in total are launched). Yet, when each sub-kernel is launched, the threads are exposed to blockIdx values ranging from 00 to $3$, causing all sub-kernels to execute the first $4$ blocks of the kernel, which will lead to erroneous results. To address this issue, we modify a kernel’s PTX code to take in an additional offset variable in its parameters and substitute occurrences of blockIdx with the sum of the offset and the original blockIdx. This approach enables threads within a sub-kernel to identify their corresponding subset of blocks. For instance, sub-kernel $2$ can reconstruct its corresponding block indices from $4$ to $7$ by adding a block offset of $4$ to the new blockIdx. With this mechanism, the collective work done by the sub-kernels can properly resemble the original kernel execution, while offering the scheduler flexible scheduling at the block-level granularity.
@@
Preemption transformation. Tally introduces a transformation that makes kernels preemptible while preserving their functional semantics. The preemption transformation in Tally is inspired by the Persistent Thread Block (PTB) programming paradigm, which deviates from the traditional programming model where the number of thread blocks launched by a kernel is proportional to the amount of work it performs. Instead, the PTB model employs a flexible number of worker blocks to perform all the work. As an example in Figure 2(a), rather than launching $16$ thread blocks of the original kernel, the PTB model launches $2$ worker blocks to process the $16$ blocks in an iterative and dynamic manner. These workers continuously fetch and increment a global integer representing the next task index. Similar to how sub-kernels in slicing utilize block offsets to recompute their block indices from the original kernel execution, workers in PTB use the task index to reconstruct the corresponding block indices for the current iteration. For example, in the illustrated case where thread blocks are structured in a $1$-dimensional grid, a fetched task index of $8$ indicates that the worker should perform the execution of the thread block with blockIdx $8$. In the case of higher-dimensional grids, block indices can be computed from the integer task index by referring to the original kernel’s grid dimension.
@@
A critical observation from the PTB model is that the progress of the kernel execution is effectively encapsulated in the global task index. Since the execution is now transformed into an iterative approach, kernels can potentially terminate early and later be resumed from the point at which they ended, using the task index. This can be achieved by introducing an additional boolean variable to the PTB kernels as a flag for signaling preemption. At the start of each iteration, workers check this flag to detect preemption signals, allowing block-level preemption within kernels. In Tally, we automatically convert kernels into this preemptive style by wrapping the original kernel function in the device code with an outer loop using control flow logic. For instance, all return statements are transformed into branch instructions that direct the execution flow to the start of the iteration, ensuring that each worker block continues execution until all tasks are processed or a preemption signal is received. Furthermore, a synchronization call is introduced at the end of each iteration to ensure that threads within a worker block process a single block at a time, preventing potential race conditions and maintaining correctness of computation.
@@
However, since the preemption transformation replaces return statements with branching statements that synchronizes threads at the start of the iteration, it can potentially lead to threads within the same block synchronizing at different places. Such divergence in synchronization is an undefined behavior in the GPU programming model and, in practice, results in infinite kernel stalls, making such transformations unsafe. To address this issue, Tally introduces a prepositional kernel transformation pass, which we denote as the unified synchronization transformation. The algorithm of the transformation is depicted in Figure 2(b). As illustrated, this transformation modifies all synchronization and return instructions by redirecting them to a newly introduced unified synchronization point that verifies if all threads are in a return state. Threads from synchronization points loop back to their previous positions, while those returning are held until all threads align. This unified synchronization transformation eliminates divergences in thread block synchronization, and in turn, the infinite stall issue. Through this prepositional transformation stage, the preemptive transformation can be applied automatically and safely.
{minted}
[ linenos, fontsize=, xleftmargin=12pt, numbersep=5pt, ]python def launch_and_profile(kernel): candidates = get_candidate_configs(kernel) for cfg in candidates: est_turnaround = lookup_measurement(kernel, cfg) if est_turnaround is None: est_turnaround = profile(kernel, cfg) if has_launched(): return set_launch_config(kernel, candidates, bound=TUNEAROUND_LATENCY_BOUND) def scheduler(): while True: for client in sort_by_priority(clients): if client.is_high_priority(): kernel = client.fetch_next_kernel() if kernel is not None: preempt_best_effort_kernel() kernel.launch(Config::DEFAULT) if client.has_kernel_running(): break else: kernel = client.get_curr_ex_kernel() if kernel is None: kernel = client.fetch_next_kernel() cfg = lookup_launch_config(kernel) if cfg is not None: kernel.launch(cfg) else: launch_and_profile(kernel) elif kernel.running(): pass elif kernel.stopped(): kernel.resume()
Figure 3. Tally priority-aware scheduling algorithm.

### v3

@@ 3. GPU SHARING IN THE WILD
High integration and maintenance cost. Many existing GPU sharing systems are fundamentally intrusive in nature, necessitating extensive modifications to user code to be used in practice. This is particularly evident in application-level systems, where users must adapt their workloads to the specific programming interfaces proposed by these solutions. Such intrusive approaches not only burden the users, but also limit their flexibility in application development (lucid, ). Similarly, runtime-level GPU sharing solutions that are designed as extensions of DL frameworks impose a significant burden on cluster managers to implement and maintain them across diverse frameworks used by the users. Consequently, deploying these solutions in practice often involves substantial integration and maintenance costs, especially given the rapidly evolving landscape of modern DL ecosystem.
@@
On the other hand, while runtime-level GPU sharing solutions that integrate seamlessly with container orchestration systems or the GPU driver may appear non-intrusive, they often impose implicit requirements on user code to provide additional application contexts needed for making effective scheduling decisions. These include the ability to temporarily launch applications in a specialized profiling environment prior to full execution (orion2024, ), or mandating applications to implement checkpoint and restart mechanisms to support live migration (gavel2020, ; xiao2018, ). However, enforcing such programming practices in large-scale, multi-tenant clusters is often impractical due to the operational complexity it introduces (Weng2022mlaas, ).
@@
Lack of performance isolation guarantees. In production DL clusters, many inference tasks are governed by service-level agreements (SLAs), which impose strict stringent requirements on their response times, typically ranging from a few milliseconds to seconds (clipper2017, ). However, current GPU sharing solutions often fall short in providing adequate performance isolation, leading to severe performance degradation for latency-critical inference tasks when co-located with other workloads (GILMAN2021102234, ). For instance, NVIDIA MPS (nvidia-mps, ) is designed to eagerly schedule as many GPU kernels from co-located tasks as possible. While this design maximizes GPU utilization, it can cause significant queuing delays for executing kernels of the high-priority inference tasks. Our experiments (see Section 5) reveal that the tail latency of a co-located inference task under MPS can increase up to 20×20$\times$20 × compared to their standalone performance.
To effectively share GPU resources while ensuring performance isolation, a GPU sharing system must prioritize one task over another and rapidly switch between them according to their dynamic resource demands. The speed of context-switching between tasks, which we denote as the turnaround latency, is dictated by the scheduling granularity employed by the GPU sharing system. As described in Section 2, the scheduling granularity used by prior works primarily fall into four categories, from coarsest to finest: iteration-level (e.g. Salus (salus2020, )), kernel-level (e.g., TGS (tgs2023, )), block-level (e.g. EffiSha (effisha2017, )), and thread-level (e.g. REEF (reef2022, )). To illustrate the effect of scheduling granularity on the tail latency, we present a representative study in Table 1 where a latency-critical BERT (devlin-etal-2019-bert, ) inference task is co-executed with a Whisper (radford2023robust, ) model training workload on an NVIDIA A100 GPU (nvidia_a100, ). In this experiment, we empirically estimate the turnaround latency for different scheduling granularities on the Whisper model and compare it to the response time of BERT inference. Since a BERT inference request takes only $3.93$ ms, achieving strong performance isolation requires turnaround latencies to be in the sub-millisecond range. However, both iteration-level and kernel-level scheduling can only achieve turnaround latencies of $3$ s and $10$ ms, respectively. As a result, scheduling the Whisper training task at these granularities can introduce 2.5760×2.5-760$\times$2.5 - 760 × higher overhead on BERT’s inference latency. In contrast, block-level and thread-level scheduling achieve microsecond-scale turnaround latencies of $304$ µs and $38$ µs, respectively, resulting in only mere $1-7\%$ overhead on the latency.
From these results, we conclude that to achieve practical performance isolation guarantees, GPU sharing solutions must schedule resources at either the block-level or thread-level granularity. Unfortunately, existing solutions either do not meet these criteria (e.g., Salus (salus2020, ), AntMan (xiao2020, ), TGS (tgs2023, )) or prove ineffective in diverse production DL clusters due to their intrusive nature (e.g., EffiSha (effisha2017, ), Orion (orion2024, )) or lack of generalizability (e.g. REEF (reef2022, )). This brings us to the third limitation.
@@
Reliance on narrowly applicable workload characteristics. Many state-of-the-art GPU sharing systems rely on specific workload characteristics in their scheduling strategies. For instance, AntMan and TGS provide runtime-level GPU sharing support but are limited to training workloads with predictable GPU kernel execution patterns. REEF, on the other hand, offers strong performance isolation for inference workloads; however, it requires all GPU kernels to be idempotent to support thread-level kernel scheduling strategies, meaning the kernels must always produce the same output for a given input, regardless of how many times it is executed. While these systems demonstrate benefits for certain classes of applications, their underlying assumptions are often not applicable to large-scale multi-tenant DL clusters that must support a diverse range of workloads with varying characteristics. More critically, prior studies (Weng2022mlaas, ) have revealed that DL clusters are often agnostic to the specific semantics of running workloads and extracting application-level context is largely infeasible. As a result, GPU sharing solutions that depend on specific task characteristics are not viable in many DL clusters.
@@
In summary, we argue that a practical GPU sharing solution for large-scale DL clusters must address all three of the aforementioned limitations.
@@ 4. TALLY: SYSTEM DESIGN
To facilitate the practical deployment of GPU sharing in production DL clusters, we propose Tally, a novel non-intrusive GPU sharing system that provides robust performance isolation for high-priority tasks during shared execution with best-effort tasks across a diverse range of DL workloads. Tally is designed as a virtualization layer that operates transparently between applications and the GPU, intercepting device API calls and employing a task-agnostic scheduling algorithm to prioritize execution from high-priority tasks while opportunistically scheduling kernels from best-effort tasks during GPU idle cycles.
@@
Tally ensures performance isolation for high-priority tasks by scheduling kernels from best-effort tasks at the block-level granularity using two scheduling primitives: slicing and preemption. Slicing divides a large kernel into numerous smaller segments, enabling more controlled scheduling, whereas preemption allows for the interruption of active kernels to expedite resource reallocation. Importantly, we make two key observations that allow applying these two scheduling primitives in a task-agnostic and non-intrusive manner. Firstly, unlike thread-level scheduling, thread blocks within a GPU are guaranteed to operate independently, making block-level scheduling universally applicable across GPU kernels. Moreover, while thread-level scheduling can provide stronger performance isolation, we find that block-level scheduling is sufficient for most practical scenarios. Secondly, the slicing and preemption scheduling primitives can be applied on GPU kernels through a series of transformation passes on the kernel device code (PTX code in the case of NVIDIA GPUs) obtained through the interception of device API calls. Leveraging these insights, we integrate the slicing and preemption primitives to Tally’s scheduling algorithm to enable microsecond-scale turnaround latency, allowing Tally to promptly switch execution between high-priority and best-effort tasks and maintain the millisecond-scale latency requirements in many latency-critical tasks.
Figure 2. Tally architecture.
@@
Figure 3. Kernel transformations: (a) slicing and preemption transformations (b) unified synchronization transformation.
Figure 2 shows the high-level design of Tally, which follows a client-server architecture. On the application side, Tally’s client library intercepts device code and API calls initiated by client processes and forwards them to the Tally server, which is responsible for managing the actual device execution and resource scheduling. The objective of the Tally server is to facilitate priority-enforced concurrent execution of one high-priority task alongside multiple best-effort tasks. To achieve this, the Tally server incorporates three key components: a (1) kernel transformer, which leverages intercepted device code to convert kernels into slicing and preemptive-style kernels, a (2) transparent profiler, which profiles the performance of the transformed kernels under various launch configurations at runtime and estimates their turnaround latencies, and finally a (3) priority-aware scheduler, which integrates capabilities of the aforementioned modules and strategically schedules GPU execution in a priority-enforced manner. In the following sections, we provide details regarding each of these components.
@@ 4.1. Kernel Transformations
Slicing transformation. Slicing is a technique that facilitates fine-grained control over kernel execution by partitioning the thread blocks of a kernel into multiple sub-kernels. Figure 3(a) illustrates this concept, where a kernel originally composed of $16$ thread blocks is divided into $4$ sub-kernels, each containing $4$ blocks. As thread blocks within a kernel are guaranteed to be independent, theoretically, this partitioning should preserve the original kernel’s functional semantics. However, naively launching the original kernel $4$ times with $4$ thread blocks each will yield incorrect results, as GPU threads rely on built-in variables such as blockIdx to determine their task assignment. Specifically, the blockIdx variable informs a thread of the index of the block it resides in within the kernel execution, allowing it to properly compute its assigned task index. In the original kernel, threads are exposed to blockIdx values ranging from 00 to $15$ (as $16$ blocks in total are launched). Yet, when each sub-kernel is launched, the threads are exposed to blockIdx values ranging from 00 to $3$, causing all sub-kernels to execute the first $4$ blocks of the kernel, which will lead to erroneous results. To address this issue, we modify a kernel’s device code to take in an additional offset variable in its parameters and substitute occurrences of blockIdx with the sum of the offset and the original blockIdx. This approach enables threads within a sub-kernel to identify their corresponding subset of blocks. For instance, sub-kernel $2$ can reconstruct its corresponding block indices from $4$ to $7$ by adding a block offset of $4$ to the new blockIdx. With this mechanism, the collective work done by the sub-kernels can properly resemble the original kernel execution, while offering the scheduler flexible scheduling at the block-level granularity.
@@
Preemption transformation. Tally introduces a transformation that makes kernels preemptible while preserving their functional semantics. The preemption transformation in Tally is inspired by the Persistent Thread Block (PTB) programming paradigm (ptb, ), which deviates from the traditional programming model where the number of thread blocks launched by a kernel is proportional to the amount of work it performs. Instead, the PTB model employs a flexible number of worker blocks to perform all the work. As an example, in Figure 3(a), rather than launching $16$ thread blocks of the original kernel, the PTB model launches $2$ worker blocks to process the $16$ blocks in an iterative and dynamic manner. These workers continuously fetch and increment a global counter which represents the next task index. Similar to how sub-kernels in slicing utilize block offsets to recompute their block indices from the original kernel execution, workers in PTB use the task index to reconstruct the corresponding block indices for the current iteration. For example, in the illustrated case where thread blocks are structured in a $1$-dimensional grid, a fetched task index of $8$ indicates that the worker should perform the execution of the thread block with blockIdx $8$. In the case of higher-dimensional grids, block indices can be computed from the integer task index by referring to the original kernel’s grid dimension.
@@
A critical observation from the PTB model is that the progress of the kernel execution is effectively encapsulated in the global task index. Since the execution is now transformed into an iterative approach, kernels can potentially terminate early and later be resumed from the point at which they ended, using the task index. This can be achieved by introducing an additional boolean variable to the PTB kernels as a flag for signaling preemption. At the start of each iteration, workers check this flag to detect preemption signals, thereby enabling block-level preemption within kernels. In Tally, we automatically convert kernels into this preemptive style by wrapping the original kernel function in the device code with an outer loop using control flow logic. For instance, all return statements are transformed into branch instructions that direct the execution flow to the start of the iteration. This ensures that each worker block will continue execution until all tasks are processed or a preemption signal is received. Furthermore, a synchronization call is introduced at the end of each iteration to make sure that threads within a worker block process a single block at a time, preventing potential race conditions and maintaining correctness of computation.
@@
Unified synchronization transformation. In the preemption transformation, return statements are substituted with branching statements that redirect threads to synchronize at the start of the iteration. However, this modification introduces a risk of threads within the same block calling synchronization at different places within the kernel. For instance, while some threads may arrive at one of the kernel’s original synchronization points, others may be redirected to the newly introduced synchronization point due to early returning. Such divergence in synchronization is an undefined behavior in the GPU programming model and, in practice, results in infinite kernel stalls. This issue makes the preemption transformation unsafe to be applied on all kernels.
To address this problem, Tally introduces a prepositional kernel transformation pass which we denote as the unified synchronization transformation. The algorithm for this transformation is depicted in Figure 3(b). As illustrated, this transformation modifies all synchronization and return instructions by redirecting them to a unified synchronization point that checks whether all threads have reached a return state. If this condition is met, all threads return simultaneously; otherwise, threads from synchronization points loop back to their previous execution positions, while those returning are held until all threads align. This modification guarantees that threads within a block return synchronously. Thus, when the kernel undergoes the preemption transformation, the risk of divergence in thread block synchronization due to early returns is eliminated, thus resolving the issue of infinite stalls. Through this prepositional transformation stage, the preemptive transformation can be applied automatically and safely to all kernels.
Figure 4. Tally’s priority-aware scheduling algorithm.